The angular acceleration of a wheel is α = 9.0t4 – 5.0t2, with α

 in radians per second-squared and t in seconds. At time t = 0, the wheel has an angular velocity of +7 rad/s and an angular position of +5 rad. Write expressions for (a) the angular velocity (rad/s) and (b) the angular position (rad) as functions of time (s).


1
Expert's answer
2022-03-20T18:50:58-0400

Given:

"\alpha = 9.0t^4 – 5.0t^2"

"\omega_0=7\:\rm rad/s"

"\theta_0=5\:\rm rad"

(a) the angular velocity

"\omega=\int \alpha dt+\omega_0=\int (9.0t^4 – 5.0t^2) dt+7\\
=1.8t^5-1.7t^3+7"

(b) the angular position

"\theta=\int \omega dt+\theta_0=\int (1.8t^5 – 1.7t^3+7) dt+5\\
=0.3t^6-0.42t^4+7t+5"